# Stream-Omni: Simultaneous Multimodal Interactions with Large Language-Vision-Speech Model

### v1

@@ 2 Related Work
For speech-oriented LMMs, existing methods rely on either continuous or discrete speech units. Methods based on continuous representations, such as Mini-Omni [8], LLaMA-Omni [9], and Freeze-Omni [26], use a speech encoder (e.g., Whisper [27]) to extract speech features, which are then projected into the LLM’s embedding space to facilitate speech understanding. These approaches often incorporate a speech decoder to generate speech responses based on LLM’s text outputs. Methods based on discrete units, such as SpeechGPT [28], Moshi [10] and GLM-4-Voice [11], employ a speech tokenizer [29, 30, 31] to convert speech into discrete units, allowing the LLM to directly understand and generate speech units, which are finally synthesized into speech using a unit-based speech decoder [32, 31]. Compared to continuous representations, discrete units can be jointly modeled with text in LLM’s context, but they often rely on more speech data for speech pre-training [33, 11].
Existing omni-modal LMMs, such as VITA-1.5 [12], MiniCPM2.6-o [7], Baichuan-Omni [13], Qwen2.5-Omni [14], use various encoders to extract the modality representations, which are then concatenated and fed into the LLM to facilitate multimodal understanding, and finally a speech decoder is employed to synthesize speech from the generated text. Such methods typically model modality alignments in a data-driven manner. In contrast, Stream-Omni models the relationships between modalities more purposefully, thereby achieving efficient and flexible modality alignments.
@@ 3.1.2 Speech Modality
In practice, to involve richer text context, Stream-Omni extends the fusion window from the aligned text token to its preceding tokens, where is the hyperparameter of window size. The alignment-based fusion is implemented via cross-attention, with the speech representations attending to the text representations, so the fused representation of speech unit is calculated as:
@@ 3.2.1 Data Construction
Stage 2: Speech-Text Alignment In this stage, the speech-text alignment is achieved by training the bottom and top speech layers using a combination of CTC loss after the bottom speech layers (refer to Eq.( 4)) and cross-entropy loss after the top speech layers. Note that, the text representations fed into the top speech layers during training (i.e., in Eq.(6)) are drawn from ground-truth transcriptions rather than LLM generated text, which aim to avoid text-speech dismatching caused by generating incorrect text, thereby enhancing the consistency of text-to-speech generation.
@@ 4.1 Benchmarks
To further assess Stream-Omni’s vision-grounded speech interaction capabilities, we construct a real-world visual-speech interaction benchmark based on the real-world VQA benchmark VisIT [56], named SpokenVisIT333https://huggingface.co/datasets/ICTNLP/SpokenVisIT. Following Fang et al. [9], the evaluation for SpokenVisIT employs the GPT model (gpt-4o version) to assign a score ranging from 1 to 5 for response. Appendix B gives the details of SpokenVisIT benchmark. Following previous works [9, 11], all speech evaluations are further divided into speech-to-text (ST) and speech-to-speech (SS) settings. For generated speech responses, we use Whisper-large-v3 [27] to transcribe the speech into text for evaluation.
@@ 4.2 Baselines
We compare Stream-Omni with vision-oriented, speech-oriented, and omni-modal LMMs of similar model scale and training data size. Vision-oriented LMM baselines include models comparable in scale to LLaVA-v1.5 [3], such as BLIP-2 [57], InstructBLIP [58], IDEFICS [59], Qwen-VL [17], Qwen-VL-Chat [17], SPHINX [19], and mPLUG-Owl2 [20]. Speech-oriented LMM baselines include TWIST [60], SpeechGPT [28], Spectron [54], Moshi [10], Freeze-Omni [26], LLaMA-Omni [9], and GLM-4-Voice [11]. Most existing omni-modal LMMs are trained on large-scale proprietary datasets. For a fair comparison, we compare Stream-Omni with VITA-1.5 [12], a text-vision-speech LMM trained on a comparable amount of data, primarily based on LLaVA [3] and LLaVA-OV [6].
4.3 Configuration
Stream-Omni is built upon the LLaMA-3.1-8B-Instruct444https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct [61], which consists of 32 Transformer layers. For vision, Stream-Omni employs the SigLIP-so400m-patch14-384555https://huggingface.co/google/siglip-so400m-patch14-384 [62] as the vision encoder. For speech, Stream-Omni incorporates the bottom speech layers with 3 Transformer layers and top speech layers with 5 Transformer layers, where all Transformer layers share the same architecture and parameter configuration as those in LLM. The speech tokenizer and flow-matching-based speech decoder are adopted from CosyVoice-300M-25Hz [31]. The vocabulary of Stream-Omni comprises 128K text tokens from LLaMA-3.1-8B-Instruct, 4096 speech units from the CosyVoice tokenizer, and a blank token . Stream-Omni is trained using 8 H800 GPUs and tested on 1 A100 GPU.
5 Results and Analyses
5.1 Visual Understanding
@@ 5.3 Vision-grounded Speech Interaction
We evaluate the vision-grounded speech interaction of Stream-Omni on the SpokenVisIT benchmark in Table 4. As the omni-modal LMMs with similar training data, Stream-Omni demonstrates superior real-world visual understanding capabilities compared to VITA-1.5. In addition, Stream-Omni supports speech generation, extending its potential for multimodal interaction. Appendix C gives specific case studies, demonstrating the advantages of Stream-Omni’s speech-text mapping in cross-modal consistency.
@@ 5.5 Effect of Alignment-based Fusion
Fusion Type For the fusion type, we compare the current cross-attention (named “Attention”) with adding aligned text representations to the input (named “Add (input)”) or each layer (named “Add (per layer)”) of the top speech layers. Results show that the attention-based approach outperforms the others, mainly due to its ability to attend to a broader context rather than merely adding a single text token.
@@ Appendix A Construction of InstructOmni
Table 1 summarizes all datasets used during training, where those marked with a superscript ‘tts’ indicate samples with synthesized speech. Overall, Stream-Omni is trained on only 23K hours of speech data, which is significantly less than the large-scale datasets used in previous methods, such as TWIST (150K hours) [60], SpeechGPT (60K hours) [28], Moshi (7M hours) [10], GLM-4-Voice (700K hours) [11], and VITA-1.5 (110K hours) [12], highlighting its clear advantage in data efficiency.
@@ Appendix B Construction of SpokenVisIT
In Sec.5.3, we construct SpokenVisIT benchmark to evaluate the vision-grounded speech interaction capability. Here, we give a detailed introduction to SpokenVisIT.
To better reflect real-world scenarios of vision-based speech interaction, we adopt the VisIT benchmark [56] as the source dataset. VisIT is a real-world visual question answering benchmark comprising 574 images and 70 types of instructions covering object recognition, visual reasoning, creative writing, and more. Unlike existing vision evaluation benchmarks that mainly use multiple-choice format, all text instructions in the VisIT benchmark are written in a colloquial style, making it particularly well-suited for speech interaction. To adapt VisIT for speech interaction, we employ text-to-speech synthesis [31] to convert each text instruction into a corresponding speech utterance, resulting in a derived benchmark named SpokenVisIT. During construction, eight math-related instructions that were unsuitable for speech interaction were removed. For the evaluation metric, following the open-ended spoken interaction evaluation protocol proposed by Fang et al. [9], we use ChatGPT (gpt-4o version) to assess the quality of model-generated responses on a 1-5 scale. The evaluation prompt includes the image caption as a reference, along with the question and the model’s answer.

### v2

@@ 2 Related Work
For speech-oriented LMMs, existing methods rely on either continuous or discrete speech units. Methods based on continuous representations, such as Mini-Omni [8], LLaMA-Omni [9], Freeze-Omni [26], SALMONN-Omni [27], and SLAM-Omni [28] use a speech encoder (e.g., Whisper [29]) to extract speech features, which are then projected into the LLM’s embedding space to facilitate speech understanding. These approaches often incorporate a speech decoder to generate speech responses based on LLM’s text outputs. Methods based on discrete units, such as SpeechGPT [30], Moshi [10] and GLM-4-Voice [11], employ a speech tokenizer [31, 32, 33] to convert speech into discrete units, allowing the LLM to directly understand and generate speech units, which are finally synthesized into speech using a unit-based speech decoder [34, 33]. Compared to continuous representations, discrete units can be jointly modeled with text in LLM’s context, but they often rely on more speech data for speech pre-training [35, 11].
Existing omni-modal LMMs, such as VITA-1.5 [12], MiniCPM2.6-o [7], Baichuan-Omni [13], Qwen2.5-Omni [14], Megrez-Omni [36], M2-Omni [37], Capybara-Omni [38], EMOVA [39], OpenOmni [40] and Omni-Emotion[41], use various encoders to extract the modality representations, which are then concatenated and fed into the LLM to facilitate multimodal understanding, and finally a speech decoder is employed to synthesize speech from the generated text. Overall, most existing omni-modal LLMs adopt concatenation-based architectures and rely primarily on data-driven approaches to model modality alignments.
In this work, we focus on the efficiency and flexibility of modality alignments, with the goal of leveraging limited tri-modal data to develop a large language-vision-speech model that can simultaneously support multimodal interactions under various modality combinations. Unlike previous approaches that rely solely on sequence-dimension concatenation for modality alignment, Stream-Omni adopts a more deliberate design by employing sequence-dimension concatenation for vision-text alignment and layer-dimension mapping for speech-text alignment, thereby achieving efficient and flexible modality alignments. As an additional advantage, the proposed layer-dimension mapping enables Stream-Omni to produce intermediate text results during speech interaction, thereby offering users a more comprehensive multimodal experience.
@@ 3.1.2 Speech Modality
In practice, to involve richer text context, Stream-Omni extends the fusion window from the aligned text token to its preceding tokens, where is the hyperparameter of window size. The alignment-based fusion is implemented via cross-attention [40], with the speech representations attending to the text representations, so the fused representation of speech unit is:
@@ 3.2.1 Data Construction
Stage 2: Speech-Text Alignment In this stage, the speech-text alignment is achieved by training the bottom and top speech layers using a combination of CTC loss after the bottom speech layers (refer to Eq.( 4)) and cross-entropy loss after the top speech layers. Note that, the text representations fed into the top speech layers during training (i.e., in Eq.(6)) are drawn from ground-truth transcriptions rather than LLM generated text, which aim to avoid text-speech dismatching [40] caused by generating incorrect text, thereby enhancing the consistency of text-to-speech generation.
@@ 4.1 Benchmarks
To further assess Stream-Omni’s vision-grounded speech interaction capabilities, we construct a real-world visual-speech interaction benchmark based on the real-world VQA benchmark VisIT [64], named SpokenVisIT333https://huggingface.co/datasets/ICTNLP/SpokenVisIT. Following Fang et al. [9], the evaluation for SpokenVisIT employs the GPT model (gpt-4o version) to assign a score ranging from 1 to 5 for response. Appendix B gives the details of SpokenVisIT benchmark. Following previous works [9, 11], all speech evaluations are further divided into speech-to-text (ST) and speech-to-speech (SS) settings. For generated speech responses, we use Whisper-large-v3444https://huggingface.co/openai/whisper-large-v3 [29] to transcribe the speech into text for evaluation.
@@ 4.2 Baselines
We compare Stream-Omni with vision-oriented, speech-oriented, and omni-modal LMMs of similar model scale and training data size. Vision-oriented LMM baselines include models comparable in scale to LLaVA-v1.5 [3], such as BLIP-2 [65], InstructBLIP [66], IDEFICS [67], Qwen-VL [17], Qwen-VL-Chat [17], SPHINX [19], and mPLUG-Owl2 [20]. Speech-oriented LMM baselines include TWIST [68], SpeechGPT [30], Spectron [62], Moshi [10], Freeze-Omni [26], LLaMA-Omni [9], and GLM-4-Voice [11]. Most existing omni-modal LMMs are trained on large-scale proprietary datasets [13, 14, 38]. For a fair comparison, we mainly compare Stream-Omni with VITA-1.5 [12], a text-vision-speech LMM trained on a comparable amount of data, primarily based on LLaVA [3] and LLaVA-OV [6]. Additionally, we also compare Stream-Omni with some methods of similar data scale to demonstrate Stream-Omni’s performance among advanced omni-modal LMMs, such AnyGPT [69], EMOVA [39] and OpenOmni [40]. Note that these models were trained using different datasets and training pipelines with Stream-Omni.
@@
4.3 Configuration
Stream-Omni is built upon the LLaMA-3.1-8B-Instruct555https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct [70], which consists of 32 Transformer layers. For vision, Stream-Omni employs the SigLIP-so400m-patch14-384666https://huggingface.co/google/siglip-so400m-patch14-384 [71] as the vision encoder. For speech, Stream-Omni incorporates the bottom speech layers with 3 Transformer layers and top speech layers with 5 Transformer layers, where all Transformer layers share the same architecture and parameter configuration as those in LLM. The speech tokenizer and flow-matching-based speech decoder are adopted from CosyVoice-300M-25Hz777https://modelscope.cn/models/iic/CosyVoice-300M-25Hz [33]. The vocabulary of Stream-Omni comprises 128K text tokens from LLaMA-3.1-8B-Instruct, 4096 speech units from the CosyVoice tokenizer, and a blank token . Stream-Omni is trained using 8 H800 GPUs and tested on 1 A100 GPU.
5 Results and Analyses
5.1 Visual Understanding
@@ 5.3 Vision-grounded Speech Interaction
Most existing benchmarks for evaluating the vision-grounded speech interaction typically use multiple-choice formats, which do not align well with real-world application scenarios. To address this, we constructed SpokenVisIT based on VisIT-Bench [64], a vision-grounded speech interaction benchmark based on real-world scenarios. We evaluate Stream-Omni on the SpokenVisIT benchmark in Table 4. As the omni-modal LMMs with similar training data, Stream-Omni demonstrates superior real-world visual understanding capabilities compared to VITA-1.5. In addition, Stream-Omni supports speech generation, extending its potential for multimodal interaction. Appendix C gives specific case studies, demonstrating the advantages of Stream-Omni’s speech-text mapping in cross-modal consistency.
@@ 5.5 Effect of Alignment-based Fusion
Fusion Type For the fusion type, we compare the current cross-attention (named “Attention”) with adding aligned text representations to the input (named “Add (input)”) or each layer (named “Add (per layer)”) of the top speech layers. Results show that the attention-based approach outperforms the others, mainly due to its ability to attend to a broader context rather than merely adding a single text token. Existing speech-oriented LMMs [9, 72] or omni-modal LMMs [40] often mix speech and text representations at the input of the speech decoder to achieve text-to-speech generation. Different with existing methods, Stream-Omni integrates the corresponding textual information into the speech representations at each layer of the speech decoder through alignment-based fusion, enabling high-quality text-to-speech generation.
@@ Appendix A Construction of InstructOmni
Table 1 summarizes the used datasets during training, where those marked with a superscript ‘tts’ indicate samples with synthesized speech. Overall, Stream-Omni is trained on only 23K hours of speech data, which is significantly less than the large-scale datasets used in previous methods, such as TWIST (150K hours) [68], SpeechGPT (60K hours) [30], Moshi (7M hours) [10], GLM-4-Voice (700K hours) [11], and VITA-1.5 (110K hours) [12], highlighting its advantage in data efficiency.
@@ Appendix B Construction of SpokenVisIT
In Sec.5.3, to align with real-world application scenarios, we construct SpokenVisIT benchmark based on VisIT-Bench [64] to evaluate the vision-grounded speech interaction capability of omni-modal LMMs. Here, we give a detailed introduction to SpokenVisIT.
To better reflect real-world scenarios of vision-based speech interaction, we adopt the VisIT-Bench [64] as the source dataset (an open-ended generation format instead of multi-choice format is much suitable for real-world scenarios). VisIT is a real-world visual question answering benchmark comprising 574 images and 70 types of instructions covering object recognition, visual reasoning, creative writing, and more. Unlike existing vision evaluation benchmarks that mainly use multiple-choice format, all text instructions in the VisIT benchmark are written in a colloquial style, making it particularly well-suited for speech interaction. To adapt VisIT for speech interaction, we employ text-to-speech synthesis [33] to convert each text instruction into a corresponding speech utterance, resulting in a derived benchmark named SpokenVisIT. During construction, eight math-related instructions that were unsuitable for speech interaction were removed. For the evaluation metric, following the open-ended spoken interaction evaluation protocol proposed by Fang et al. [9], we use ChatGPT (gpt-4o version) to assess the quality of responses on a 1-5 scale. The evaluation prompt includes the image caption as a reference, along with the question and the model’s answer.